OpenStudy (anonymous):

Solve the following system of equations. x + 4y + z = –10 3x – 3y + 6z = –21 x + 2y + 2z = –10 (1, –2, –3) (–4, –1, –2) (–3, –2, –3) (2, –2, –4) Also, if anyone would like to, I was wondering if they could help me with 2 more questions that are the same.

OpenStudy (anonymous):

I don't really get how to solve it with 3 variables know and I was wondering if anyone could help me understand it.

OpenStudy (lgbasallote):

use systems of equations on two equations first...i'll choose 1 and 3 (because they're simple) x + 4y + z = -10 x + 2y + 2z = -10 now i subtract the equations 0 + 2y - z = 0 2y = z <--take note of this..i'll use it later i sub this value of z into the first equation x + 4y + z = -10 x + 4y + 2y = -10 x + 6y = -10 x = -10 - 6y <---take note of this too now i'll sub in x and z into the second equation 3x - 3y + 6z = -21 i'll simplify it first by dividing each term by 3 x - y + 2z = -7 now i substitute (-10 - 6y) - y + 2(2y) = -7 -10 - 6y - y + 4y = -7 -3y - 10 = -7 -3y = 10 - 7 -3y = 3 y = -1 now you have a value of y...can you look for z and x now?

OpenStudy (anonymous):

I'm sort of confused. When you added them at beginning wouldn't one of the x's have to be negative to cancel out? Also, wouldn't that be -20?

OpenStudy (lgbasallote):

i didnt add..i subtracted x - x = 0 4y - 2y = 2y z - 2z = -z -10 -(-10) = -10 + 10 = 0

OpenStudy (anonymous):

Oh I didn't know that. I would always multiply one side so one of the values would cancel out. K thx.

OpenStudy (lgbasallote):

well the moment i saw some single variables i knew i had to use them to make life simple hehe

OpenStudy (anonymous):

Wouldn't x=-4 and z=2?

OpenStudy (lgbasallote):

hmm...remember look at the equations i pointed for you to remember 2y = z x = -10 - 6y

OpenStudy (anonymous):

z=-2*

OpenStudy (lgbasallote):

now that's right ^_^

OpenStudy (anonymous):

Lol sry was a typo. DId it right but wrote it wrong

OpenStudy (anonymous):

Tyvm lgba :)

OpenStudy (lgbasallote):

my pleasure ^_^ \[\Huge \color{maroon}{\mathtt{\text{<tips hat>}}}\]

OpenStudy (anonymous):

The following system is solved by substitution. Which expression is substituted for z in the second and third equations? x + 5y – z = 6 3x – y – 2z = 4 3x – 2y + 3z = 13 x + 5y + 6 x + 5y – 6 <===== My answer –x – 5y + 6 –x – 5y – 6 Am I right?

OpenStudy (anonymous):

lol :) could you just check this real quick?

OpenStudy (lgbasallote):

yup that's right

OpenStudy (anonymous):

K thx. and lol got one more question I'll try first on my virtual quiz.

OpenStudy (anonymous):

If I need any help, I'll post it here :)

OpenStudy (lgbasallote):

well if you have a question abt it..just post as a new question..i gtg now sorry

OpenStudy (anonymous):

K no problem, and thx for all the help :)